OpenStudy (anonymous):

**Correct Me [Work Is Shown] Francine installed a new pool for $11,320 using a 12-month deferred payment plan with an interest rate of 20.67%. What is the balance after the deferment period if payments of $436 are made each month? $6,088.00 $8,662.75 $11,320.00 $13,894.75<------ ** MY WORK *** A=11.320(1+0.2067/12)^(12*1) = 13894.75 ?

OpenStudy (amistre64):

this is just a normal run of compounding with payments

OpenStudy (amistre64):

\[B_{12}=11320(k)^{12}-436\frac{1-(k)^{12}}{1-(k)}\] where k = 1+.2067/12

OpenStudy (amistre64):

paying off the balance over the defered period would not result in a higher balance to be paid

OpenStudy (amistre64):

so essentialy you found the balance that is componded over the period, but did not adjust the payments from it

OpenStudy (anonymous):

so k=1.017225

OpenStudy (amistre64):

yes, which you can rnd to 4 decies most likeley

OpenStudy (anonymous):

b\[B_{12} =11320(1.1017225)^{12} -436^{1-1.017225}^{12} divided by 1-(1.017225)\]

OpenStudy (anonymous):

B=11320(1.017225)^(12)-436^1-(1.017225)^(12)/1-(1.017225)

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

B=11320(1.017225)^(12)-436(1-(1.017225)^(12))/(1-(1.017225)) i dont get an exact option with that, but it should get you close enough